# Danish Fungi 2024 DF20 Danish Fungi 2020

### v2

@@ Abstract
We introduce a new, challenging benchmark and a dataset, FungiTastic, based on fungal records continuously collected over a twenty-year span. The dataset is labeled and curated by experts and consists of about 350k multimodal observations of 5k fine-grained categories (species). The fungi observations include photographs and additional data, e.g., meteorological and climatic data, satellite images, and body part segmentation masks. FungiTastic is one of the few benchmarks that include a test set with DNA-sequenced ground truth of unprecedented label reliability. The benchmark is designed to support (i) standard closed-set classification, (ii) open-set classification, (iii) multi-modal classification, (iv) few-shot learning, (v) domain shift, and many more. We provide tailored baselines for many use cases, a multitude of ready-to-use pre-trained models on HuggingFace, and a framework for model training. A comprehensive documentation describing the dataset features and the baselines are available at GitHub and Kaggle.
@@ 1 Introduction
Biological problems provide a natural, challenging setting for benchmarking image classification methods [40, 51, 52]. Consider the following aspects inherently present in biological data. The species distribution is typically seasonal and influenced by external factors such as recent precipitation levels. Species categorization is fine-grained, with high intra-class and inter-class variance. The distribution is often long-tailed; for rare species, only a few samples are available. New species are being discovered, raising the need for the “unknown” class option. Commonly, the set of classes has a hierarchical structure, and different misclassifications may have very different costs. Think of mistaking a poisonous mushroom for an edible one, potentially lethal, and an edible mushroom for a poisonous one, which at worst means returning home with an empty basket. Similarly, needlessly administering anti-venom after making a wrong decision about a harmless snake bite may be unpleasant, but its consequences are incomparable to not acting after a venomous bite.
@@
The properties of biological data listed above enable testing of, e.g., both open-set and closed-set categorization, robustness to prior and appearance domain shift, performance with limited training data, and dealing with non-standard losses. In contrast, most common benchmarks [10, 53] operate under the independent and identically distributed (i.i.d.) assumption, which is made valid by shuffling data and randomly splitting it for training and evaluation. In real-world applications, i.i.d data are rare since training data are collected well before deployment and everything changes over time [55].
@@
Data sources play an important role in benchmarking [19, 21]. In the age of LLMs and VLMs trained on possibly the entirety of the internet at a certain point in time, access to new, “unseen” data is crucial to ensure methods are tested on data not indirectly “seen”, without knowing. Conveniently, many domains in nature are of interest to experts and the general public, who both provide a continuous stream of new and annotated data [39, 49]. The general public’s involvement introduces the problem of noisy training data; evaluating robustness to this phenomenon is also of practical importance.
Classification of data originating in nature, including images of birds [3, 53], plants [15, 18], snakes [6, 36], fungi [40, 52], and insects [17, 32] has widely been used for general benchmarking of machine learning algorithms, not just for fine-grained visual categorization. Most of the commonly used datasets are small for current standards; the number of classes is also limited. The performance is often saturated, reaching total accuracy between 85–95 %; see the rightmost column of Tab. 1. Typically, the datasets are solely image-based and focused on traditional image classification; few offer basic metadata attributes. Moreover, many popular datasets suffer from specific problems, e.g., regional, racial, and gender biases [48] and errors in labels [4, 50].
These limitations are particularly problematic for deployment, where machine learning models are expected to work with changing distributions and variable categories. As a result, the necessity of more comprehensive datasets that incorporate diverse modalities, reduce bias, and support more advanced learning paradigms, including multimodal, open-set, and few-shot classification, are becoming increasingly clear to better reflect the complexity of real-world data.
@@
Modalities SOTA
Dataset + citations (2022-24) Classes Training Test I M S Tasks Accuracy
Oxford-IIIT Pets [34] 1,060 37 1,846 3,669 C 97.1 [13]
FGVC Aircraft [31] 1,190 102 6,732 3,468 C 95.4 [2]
Stanford Dogs [26] 680 120 12,000 8580 C 97.3 [2]
Stanford Cars [28] 2,060 196 8,144 8,041 C 97.1 [29]
CUB-200-2011 [53] 1,910 200 5,994 5,794 C 93.1 [7]
PlantNet300k [16] 30 1,081 243,916 31,112 C 92.4 [16]
ImageNet-1k [10] 21,200 1,000 1,281,167 100,000 C, FS 92.4 [12]
iNaturalist [52] 727 5,089 579,184 95,986 C, FS 93.8 [46]
ImageNet-21k [42] 456 21,841 14,197,122 - C, FS 88.3 [46]
Insect-1M [32] 8 34,212 1,017,036 - C, MM
Species196 [20] 5 196 9,676 9,580 C, MM 88.7 [20]
NABirds [50] 283 555 48,562 - C, FS, MM 93.0 [11]
(our) DF20 [40] 42 1,604 266,344 29,594 C 80.5 [40]
(our) DF20–Mini [40] 42 182 32,753 3,640 C 75.9 [40]
FungiTastic 2,829 433,701 91,832 $\forall$ 75.3
FungiTastic–Mini 215 46,842 10,738 $\forall$ 74.8
In the paper, we introduce FungiTastic, a comprehensive multi-modal dataset of fungi observations [^1] A set of photographs and additional metadata describing one particular fungi specimen and surrounding environment. Usually, each photograph focuses on a different organ. For example observation, see Figure1 which takes advantage of the favorable properties of natural data discussed above. The fungi observations include photographs, satellite images, meteorological data, segmentation masks, and location-related metadata. The location metadata enriches the observations with attributes such as the timestamp, GPS location, and information about the substrate and habitat. By incorporating various modalities, the dataset supports a robust benchmark for multi-modal classification, enabling the development and evaluation of sophisticated machine-learning models under realistic and dynamic conditions. The key contributions of the proposed FungiTastic benchmark are:
  • It includes diverse data types, such as photographs, satellite images, bioclimatic time-series data, segmentation masks, and contextual metadata (e.g., timestamp, camera metadata, location, substrate, and habitat), offering a rich, multimodal benchmark.
  • It addresses real-world challenges such as domain shifts, open-set problems, and few-shot classification, providing a realistic benchmark for developing robust machine learning models.
  • The proposed benchmarks allow for addressing fundamental problems such as novel-class discovery, few-shot classification, and evaluation with non-standard cost functions.
User provided photographs Body part masks Satellite Climatic data
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Externally crawled metadata User provided observation metadata
 Toxicity: Edible Elevation: 28.5m Substrate: Soil Habitat: Nat. grassland
Bio zone: Atlantic Landcover: Grasslands Location: 56.84, 9.01 Date: 2023-10-13
Figure 1: A FungiTastic observation includes one or more photographs of an observed specimen and usually also a microscopic image of its spores. For a subset ($\sim$70k photographs), we provide body part segmentation masks. The user-provided observation metadata is available for virtually all observations as well as additional crawled metadata. Geospatial data is available for all observations, DNA sequencing for a subset. Last but not least, we provide climatic time-series data.
@@ 2 The FungiTastic Benchmark
FungiTastic is built from fungi observations submitted to the Atlas of Danish Fungi before the end of 2023, which were labeled by taxon experts on a species level. In total, more than 350k observations consisting of 650k photographs collected over 20 years are used. Apart from the photographs, each observation includes additional observation data (see Figure 1) ranging from satellite images, meteorological data, and tabular metadata (e.g., timestamp, GPS location, and information about the substrate and habitat) to segmentation masks and toxicity status. The vast majority of observations got most of the attributes annotated. For more details about the attribute description and its acquisition process, see Subsection 2.1.
Since the data comes from a long-term conservation project, its seasonality and naturally shifting distribution make it suitable for time-based splitting. In this so-called temporal division, all data collected before a certain point is used for training, while data collected after that point is reserved for evaluation. Specifically, in our case, the train datasets consist of all observations up to the end of 2021, and the val and test datasets include all observations from 2022 and 2023, respectively.
The FungiTastic benchmark is designed to support not just standard closed-set classification but also (i) open-set classification, (ii) few-shot learning, (iii) multi-modal classification, (iv) domain shift, and many more. For the most common task, we provide a short description of curated subsets below and additional quantitative information about all the subsets in the Appendix in Table 6.
@@
FungiTastic is a general subset that includes most of the observations (around 340k) and species (i.e., 2,829) alongside all the additional observation data except for the body part masks. As a bonus, we include a DNA-based test set of 725 species and 2,041 observations. The FungiTastic comes along with dedicated validation and test sets specifically designed for closed-set and open-set scenarios. While the closed-set validation and test sets only include species present in the training set, the open-set also includes observations with species that were observed only after 2022 and 2023, i.e., species not available in the training set. All the species (i.e., categories) with no examples in the training set are labeled as "unknown".
@@
FungiTastic–Mini is a compact and challenging subset of the FungiTastic dataset designed primarily for prototyping and consisting of all observations belonging to 6 hand-picked genera (e.g., Russula, Boletus, Amanita, Clitocybe, Agaricus, and Mycena) [^2] These genera produce fruiting bodies of the toadstool type, which include many visually similar species and are of significant interest to humans due to their common use in gastronomy.. This subset comprises 46,842 images (25,786 observations) of 215 species, greatly reducing the computational requirements for training.
@@
FungiTastic–FS subset, FS for few-shot, is formed by species with less than 5 observations in the training set, which were removed from the main (FungiTastic) dataset. The subset contains 4,293 observations encompassing 7,819 images of a total of 2,427 species. As in the FungiTastic data, the split into validation and testing is done according to the year of acquisition.
@@ 2.1 Additional Observation Data
This section provides a comprehensive overview of the accompanying data available alongside the user-submitted photographs. For each subset, we describe the data itself and, if needed, its acquisition process as well. Below, we describe (i) tabular metadata, which include key environmental attributes and taxonomic information accompanying nearly all image observations, (ii) remote sensing data at fine-resolution geospatial scale for each observation site, (iii) meteorological data, which provides long-term climate variables, and (iv) body part segmentation masks that delineate specific morphological features of fungi fruiting bodies, such as caps, gills, pores, rings, and stems. All those metadata are integral to advancing research combining visual, textual, environmental, and taxonomic information.
Multi-band remote sensing data such as satellite images offer detailed and globally consistent environmental information at a fine resolution, making it a valuable resource for species categorization (i.e., identification) [44, 35]. To allow testing the potential of such data and to facilitate easy use of geospatial data, we provide multi-band (e.g., R, G, B, NIR, and elevation) satellite patches with 128$\times$128 pixel resolution (10m spatial resolution per pixel for R, G, B, and NIR, and 30m for elevation), centered on observation location. The data were extracted from rasters publicly available at Ecodatacube and ASTER Global Digital Elevation Model. While the extraction of the elevation data is straightforward, the raw pixel values in the Sentinel-2A rasters available on Ecodatacube might include extreme values. Therefore, we had to post-process raw data further to be in a standardized and expected form (i.e., an image with [0, 255] values). First, we clipped the values at 10,000. Next, the values were rescaled to a [0, 1] range and adjusted with a gamma correction factor of 2.5 (i.e., the values were raised to the power of 1/2.5). At last, the values were rounded and rescaled to [0, 255].
Figure 2: Satellite images. RGB images with a 128$\times$128 resolution extracted from Sentinel2 data.
@@
Body part segmentation masks of fungi fruiting body is essential for accurate identification and classification [9]. These morphological features provide crucial taxonomic information distinguishing some visually similar species. Therefore, we provide human-verified instance segmentation masks for all photographs in the FungiTastic–Mini (FungiTastic–M). We consider various semantic categories such as caps, gills, pores, rings, stems, etc. These annotations are expected to drive advances in interpretable recognition methods [43] and evaluation [22], with masks also enabling instance segmentation for separate foreground and background modeling [5]. All segmentation mask annotations were semi-automatically generated in CVAT using the Segment Anything Model [27].
@@
Figure 3: Fruiting body part segmentation. We consider cap, gills, stem, pores, and ring.
@@
Location-related metadata is provided for approximately 99.9% of the observations and describes the location, time, taxonomy and toxicity of the specimen, surrounding environment, and capturing device. See Table 2 for a detailed description of all available location-related metadata. While part of the metadata is usually provided by citizen scientists [^3] A member of the public who actively participates in scientific research and data collection, contributing valuable information and observations to support professional scientists. some attributes (e.g., elevation, land cover, and biogeographical) are crawled externally; all with a great potential to improve the classification accuracy and enable research on combining visual data with metadata.
@@
Table 2: List of available location-related metadata. For a predominant number of observations, we provide a comprehensive set of data describing the surroundings or the specimen itself. Using such metadata for species identification allows to improve accuracy; see [11, 40].
@@
Date of observation Date when the specimen was observed in a format yyyy-mm-dd. Besides, we provide three additional columns with pre-extracted year, month, and day values.
EXIF Camera device attributes extracted from the image, e.g., metering mode, color space, device type, exposure time, and shutter speed.
@@
Substrate The natural substance on which the specimen lives. A total of 32 values such as Bark, Soil, Stone, etc.
@@
Toxicity status Whether is the species toxic or no. Since non-edible species can cause serious health issues as well, we include them as poisonous.
Location Location data are provided in various formats, all upscaled from decimal GPS coordinates. Besides the latitude and longitude, we also provide administrative divisions for regions, districts, and countries.
@@
Elevation Standardized elevation value, i.e., height above the sea level.
Landcover A land cover classification encodings including values such as savanna, barren, etc. Taken from MODIS Terra+Aqua [14].
Meteorological data and other climatic variables are vital assets for species identification and distribution modeling [1, 23]. In light of that, we provide 20 years of historical time-series monthly values of mean, minimum, and maximum temperature and total precipitation for all observations (see Figure 4 for example data). For each observation site, 20 years of data was extracted; for instance, an observation from the year 2000 includes data from 1980 to 2000. However, as the available climatic rasters only extend up to the year 2020, observations from 2020 to 2024 have missing values for those years not covered by existing data. In addition, we also provide 19 bioclimatic variables (related to temperature, seasonality, etc.) averaged over the period from 1981 to 2010. All data were extracted from climatic rasters available at CHELSA [25, 24].
Figure 4: Sample of available temperature data. First 5 years (out of 20 available years) of average, maximum, and minimum monthly temperatures for a randomly selected location.

### v3

@@ Abstract
We introduce a new, challenging benchmark and a dataset, FungiTastic, based on fungal records continuously collected over a twenty-year span. The dataset is labelled and curated by experts and consists of about 350k multimodal observations of 6k fine-grained categories (species). The fungi observations include photographs and additional data, e.g., meteorological and climatic data, satellite images, and body part segmentation masks. FungiTastic is one of the few benchmarks that include a test set with DNA-sequenced ground truth of unprecedented label reliability. The benchmark is designed to support (i) standard closed-set classification, (ii) open-set classification, (iii) multi-modal classification, (iv) few-shot learning, (v) domain shift, and many more. We provide tailored baselines for many use cases, a multitude of ready-to-use pre-trained models on HuggingFace, and a framework for model training. The documentation and the baselines are available at GitHub and Kaggle.
Figure 1: A FungiTastic observation includes one or more photos of an observed specimen with expert-verified taxon labels (some DNA sequenced) and occasionally also a microscopic image of its spores. Textual captions, observation metadata, geospatial data, and climatic time-series data are available for virtually all observations. For a subset ($\sim$70k photos), we provide body part segmentation masks.
@@ 1 Introduction
Biological problems provide a natural, challenging setting for benchmarking image classification methods [49, 66, 65, 56]. Consider the following aspects inherently present in biological data. The species distribution is typically seasonal and constantly evolving under the influence of external factors such as precipitation levels, temperature, and loss of habitat, exhibiting constant domain shifts. Species categorization is fine-grained, with high intra-class and low inter-class variance. The distribution is often long-tailed; only a few samples are available for rare species (few-shot learning). New species are being discovered, raising the need for the “unknown” class option (i.e., open-set recognition). Commonly, the set of classes has a hierarchical structure, and different misclassifications may have very different costs (i.e., non-standard losses). Think of mistaking a poisonous mushroom for an edible, potentially lethal, and an edible mushroom for a poisonous one, which at worst means an empty basket. Similarly, needlessly administering anti-venom after making a wrong decision about a harmless snake bite may be unpleasant, but its consequences are incomparable to not acting after a venomous bite.
@@
Common benchmarks [15, 68, 43, 65] generate independent and identically distributed (i.i.d.) data by shuffling and randomly splitting it for training and evaluation. In real-world applications, i.i.d data are rare since training data are collected well before deployment and everything changes over time [70]. Moreover, they fail to address the above-mentioned aspects important in many instance of ML system deployment: robustness to distribution and domain shifts, ability to detect classes not represented in the training set, limited training data, and dealing with non-standard losses.
@@
For benchmarking, it is crucial to ensure that methods are tested on data not indirectly “seen”, without knowing  [24, 27], especially given the huge dataset used for training LLMs or VLMs and possibly covering the entirety of the internet at a certain point in time. Conveniently, many domains in nature are of interest to experts and the general public, who provide a continuous stream of new and annotated data [48, 61]. The public’s involvement introduces the problem of noisy training data; evaluating the robustness of this phenomenon is also of practical importance.
In the paper, we introduce FungiTastic, a multi-modal dataset and benchmark based on fungi observations [^1] A set of photographs and additional metadata describing one particular fungi specimen and surrounding environment. Usually, each photograph focuses on a different organ. For an example observation, see Figure 1, which takes advantage of the favorable properties of natural data discussed above and shown in Figure 1. The fungi observations include photographs, satellite images, meteorological data, segmentation masks, textual captions, and location-related metadata. The location metadata enriches the observations with attributes such as the timestamp, GPS location, and information about the substrate and habitat.
By incorporating various modalities, the dataset supports a robust benchmark for multi-modal classification, enabling the development and evaluation of sophisticated machine-learning models under realistic and dynamic conditions.
The key contributions of the FungiTastic benchmark are:
  • It addresses real-world challenges such as domain shifts, open-set problems, and few-shot classification, providing a realistic benchmark for developing robust ML models.
  • The proposed benchmarks allow for addressing fundamental problems beyond standard image classification, such as novel-class discovery, few-shot classification, and evaluation with non-standard cost functions.
  • It includes diverse data types, such as photographs, satellite images, bioclimatic time-series data, segmentation masks, contextual metadata (e.g., timestamp, camera metadata, location, substrate, and habitat), and image captions, offering a rich, multimodal benchmark.
2 Related Work
Classification of data originating in nature, including images of birds [6, 68], plants [21, 23], snakes [9, 47], fungi [49, 65], and insects [22, 44] has been widely used to benchmark machine learning algorithms, not just fine-grained visual categorization. The datasets were instrumental in focusing on fine-grained recognition and attracting attention to challenging natural problems.
However, the datasets are typically artificially sampled, solely image-based, and focused on traditional image classification. Most commonly used datasets are small by modern standards, with a limited number of categories, which restricts their usefulness for large-scale and highly diverse applications. Though performance being often saturated, reaching an accuracy of 85–95 % (rightmost column of Tab. 1), these datasets are still widely used in the community and have reached thousands of citations in the past few years. Many popular datasets also suffer from specific limitations that compromise their generalizability and robustness. Common issues include:
  • Lack of Multi-Modal Data: Available datasets are predominantly image-based, with few offering auxiliary metadata like geographic or temporal context, which is essential for real-world applications where distribution changes and context is important.
  • Biases in Data Representation: Many datasets exhibit regional and other biases [60], which can lead to biased models that do not perform well across different populations or environments. This lack of diversity can severely limit the usability of models trained on these datasets for global applications.
  • Single task focus: While current ML applications require adaptability to tasks such as open-set classification, few-shot learning, and out-of-distribution detection, many of these datasets were not designed with these tasks in mind, limiting their usefulness for modern benchmarking.
  • Labeling Errors and Quality Control: Label errors are prevalent in widely-used datasets [7, 64]. Mislabeling, especially in fine-grained categories, can reduce the reliability of these datasets as benchmarks and reduce the model’s ability to learn fine distinctions.
@@ 2 Related Work
Modals. SOTA
Dataset Classes Images I M S Tasks Accuracy
Oxford-Pets [46] 37  5k C 97.1 [19]
FGVC Aircraft [43] 102  10k C 95.4 [5]
Stanford Dogs [34] 120  20k C 97.3 [5]
Stanford Cars [36] 196  16k C 97.1 [38]
Species196 [26] 196 20k C///M2 88.7 [26]
CUB-200-2011 [68] 200  12k C 93.1 [11]
NABirds [64] 555 49k C///F///M2 93.0 [16]
PlantNet300k [21] 1,081  275k C 92.4 [21]
DanishFungi2020 [49] 1,604 296k  C///M2 80.5 [49]
ImageNet-1k [15] 1,000 1.4m C///FS 92.4 [17]
TaxaBench-8k [56] 2225 9k  C///M2 37.5  [56]
iNaturalist [65] 5,089 675k C///FS 93.8 [58]
ImageNet-21k [52] 21,841 14m C///FS 88.3 [58]
Insect-1M [44] 34,212 1m  C///M2
(our) FungiTastic 2,829 620k $\forall$ 75.3
(our) FungiTastic–Mini 215 68k $\forall$ 74.8
@@ 3 The FungiTastic Benchmark
FungiTastic is built from fungi observations submitted to the Atlas of Danish Fungi before the end of 2023, which were labeled by taxon experts on a species level. In total, more than 350k observations consisting of 630k photographs collected over 20 years are used. Apart from the photographs, each observation includes additional observation data (see Figure 1) ranging from satellite images, meteorological data, and tabular metadata (e.g., timestamp, GPS location, and information about the substrate and habitat) to segmentation masks and toxicity status. The vast majority of observations got all of the attributes annotated. For details about the attribute description and its acquisition process, see Subsection 3.1. Since the data comes from a long-term conservation project, its seasonality and naturally shifting distribution make it suitable for time-based splitting. In this so-called temporal division, all data collected up to the end of 2021 is used for training, while data from 2022 and 2023 is reserved for validation and testing, respectively.
The FungiTastic benchmark is designed to go beyond standard closed-set classification and support a wide range of challenging machine learning tasks, including (i) open-set classification, (ii) few-shot learning, (iii) multi-modal learning, and (iv) domain shift evaluation. To facilitate these tasks, we provide several curated subsets, each tailored for specific experimental setups. A general overview of these subsets is provided below, with detailed statistics and further information in the Appendix (see Table 9).
@@
FungiTastic is a general subset that includes around 346k observations of 4,507 species accompanied by a wide set of additional observation data. The FungiTastic has dedicated validation and test sets specifically designed for closed-set and open-set scenarios. While the closed-set validation and test sets only include species present in the training set, the open-set also includes observations with species observed only after 2022 (validation) and 2023 (test), i.e., species not available in the training set. All the species with no examples in the training set are labeled as "unknown". Additionally, we include a DNA-based test set of 725 species and 2,041 observations.
@@
FungiTastic–Mini (FungiTastic–M) is a compact and challenging subset of the FungiTastic dataset designed primarily for prototyping and consisting of all observations belonging to 6 hand-picked genera (e.g., Russula, Boletus, Amanita, Clitocybe, Agaricus, and Mycena) [^2] These genera produce fruiting bodies of the toadstool type, which include many visually similar species and are of significant interest to humans due to their common use in gastronomy.. This subset comprises 67,848 images (36,287 observations) of 253 species, greatly reducing the computational requirements for training. Exclusively, we include body part mask annotations.
@@
FungiTastic–FS subset, FS for few-shot, is formed by species with less than 5 observations in the training set, which were removed from the main (FungiTastic) dataset. The subset contains 6,391 observations encompassing 12,015 images of a total of 2,427 species. As in the FungiTastic data, the split into validation and testing is done according to the year of acquisition.
@@ 3.1 Additional Observation Data
This section provides an overview of the accompanying data available for virtually all user-submitted observations. For each type, we describe the data itself and, if needed, its acquisition process as well. Below, we describe (i) tabular metadata, which includes key environmental attributes and taxonomic information for nearly all observations, (ii) remote sensing data at fine-resolution geospatial scale for each observation site, (iii) meteorological data, which provides long-term climate variables, (iv) body part segmentation masks that delineate specific morphological features of fungi fruiting bodies, such as caps, gills, pores, rings, and stems, and (v) image captions. All that metadata is integral to advancing research combining visual, textual, environmental, and taxonomic information.
@@
Body part segmentation masks of fungi fruiting bodies are essential for accurate identification and classification [13]. These morphological features provide crucial taxonomic information distinguishing some visually similar species. Therefore, we provide instance segmentation masks for all photographs in the FungiTastic–M. We consider various semantic categories such as cap, gills, pores, ring, stem, etc. These annotations (see Figure 2) are expected to drive advances in interpretable recognition methods [53] and evaluation [29], with masks also enabling instance segmentation for separate foreground and background modeling [8]. All segmentation mask annotations were semi-automatically generated in CVAT using the Segment Anything Model [35] and human supervision, i.e., annotators fixed all wrong masks.
@@
Figure 2: FungiTastic body part segmentation. We consider five different categories, e.g., the cap, gills, stem, pores, and the ring.
Multi-band remote sensing data offer detailed and globally consistent environmental information at a fine resolution, making it a valuable resource for species categorization (i.e., identification) [54] and species distribution modeling [10, 50]. To allow testing the potential of such data and to facilitate easy use of geospatial data, we provide multi-band (e.g., R, G, B, NIR, elevation, and landcover) satellite patches with 64$\times$64 pixel resolution at 10m spatial resolution per pixel for (elevation and landcover are re-projected from 30m), centered on observation location. The data were extracted from rasters publicly available at Ecodatacube, ASTER, and ESA WorldCover. The data are available in the form of torch tensors in a shape of [$6\times 64\times 64$].
Figure 3: Satellite RGB images with 64$\times$64 resolution extracted from Sentinel-2A rasters available at Ecodatacube.
Meteorological data and other climatic variables are vital assets for species identification and distribution modeling [4, 30]. In light of that, we provide 20 years of historical time-series monthly values of mean, min., and max. temperature and total precipitation for all observations (see Figure 8 in Appendix for example data). For each observation site, 20 years of data was extracted; for instance, an observation from 2000 includes data from 1980 to 2000. However, as the available climatic rasters only extend up to the year 2020, observations from 2020 to 2024 have missing values for those years not covered by existing data. In addition, we provide 19 bioclimatic variables (e.g., temp., seasonality, etc.) averaged over the period from 1981 to 2010. All data were extracted from CHELSA [33, 32].
Image captions. Recent advances in VLMs [37, 1, 2] have demonstrated strong performance across tasks such as image reasoning [1] and captioning [37] and shown that VLMs can effectively understand and reason about fine-grained details within images [39]. Building on these insights, we provide text descriptions for most photographs using the state-of-the-art open-source Malmo-7B VLM model [14]. We generate baseline captions (see Figure 4 and Figure 9, App.) with a prompt specifically designed to emphasize visual characteristics relevant to fungi identification, while avoiding unnecessary or potentially misleading details. The following prompt was used to guide the caption generation:
“Describe the visual features of the fungi, such as their colour, shape, texture, and relative size. Focus on the fungi and their parts. Provide a detailed description of the visual features, but avoid speculations.”
Refer to caption Its stem is thick and light brown, with a hint of green at the base. The smaller mushroom on the right has a similar light brown cap, but its rim is more pronounced and has a white, almost translucent appearance. This gives it a delicate, lacy look. The stem of this mushroom is thinner and lighter in color compared to ……
Figure 4: Image caption sample. For each photograph, we use a Malmo-7B [14] VLM to produce a realistic image caption with an exhaustive text description.
@@
Location-related metadata is provided for approximately 99.9% of the observations and describes the location, time, taxonomy, and toxicity of the specimen, surrounding environment, and capturing device. See Table 2 for a detailed description of all available location-related metadata. While part of the metadata is usually provided by citizen scientists [^3] A member of the public who actively participates in data collection, contributing valuable information to support professional scientists., some attributes (e.g., elevation, land cover, and biogeographical) are crawled externally; all with potential to improve the classification accuracy and enable research on combining visual data with metadata.
@@
Table 2: List of available location-related metadata. For virtually all observations (>99.9%), we provide data describing the surroundings or the specimen. Using such data for species identification allows to improve accuracy; see [16, 49].
@@
Observation date Date when the specimen was observed in yyyy-mm-dd format. Besides, three additional columns with pre-extracted year, month, and day values are provided.
EXIF Camera attributes extracted from the image, e.g., metering mode, color space, device type, exposure, etc.
@@
Substrate The natural substance on which the specimen lives. Selected from 32 values such as Bark, Soil, Stone, etc.
@@
Toxicity status Whether the species is poisonous or not as a binary value. Since non-edible species can cause serious health issues as well, we label them as poisonous.
Location Latitude + longitude and coarser administrative divisions into regions, districts, and countries.
@@
Elevation Standardized height above the sea level.
Landcover Land cover classification code with values like savanna, barren, etc. Taken from MODIS Terra+Aqua [20].
Figure 5: Class distribution shift on the FungiTastic–M dataset.The long-term data acquisition captures a phenomenon related to natural changes in species presence, i.e., class prior shift. Sorted in descending order based on their occurrence in the training set. The training set includes data from 2021 and before (215 species), the validation set from 2022 (196 species), and the test set from 2023 (193 species).